You create a lab to investigate the concept of wor
Answer #1

a) Here, we see that at intial, mass is hanging and force sensor is not in motion at that instant,

so, F = mg = 0.2*9.81 = 1.962 N

But if we consider that mass m is moving with constant acceleration a .If we draw the FBD for both mass,m and Cart+string (M) and let T be tension in string => T=m(a+g) and T= Ma => a = (mg/M-m) = 0.2*9.81 = 1.962 m/s

And F = (m+M)*a = 1.4*1.962 = 2.7468 N = 2.7 N (approx)

b) Now cart+sensor are in motion.And work-energy say, change in KE = Workdone

mgh = 1/2 m'v2